OpenStudy (joftheworld):

hey can you help me plot three points from a few linear equations ? its difficult because I need to find whole numbers for the points...

OpenStudy (triciaal):

do you have the equations?

OpenStudy (joftheworld):

yes heres one.. -3y=5x-7

OpenStudy (triciaal):

linear equation y = mx + b

OpenStudy (triciaal):

why do you need whole numbers?

OpenStudy (joftheworld):

because its a online test.. and the graph wont let me plot the point in between numbers..

OpenStudy (triciaal):

did you try (0, 7)

OpenStudy (joftheworld):

I need three points..

OpenStudy (triciaal):

what is the original question?

OpenStudy (joftheworld):

-3y=5x-7

OpenStudy (triciaal):

what is the original question? with the directions

OpenStudy (triciaal):

are all the linear equations in the same question to be solved simultaneously?

OpenStudy (joftheworld):

it just says graph the equation by plotting 3 points

OpenStudy (triciaal):

(-4, 9)

OpenStudy (joftheworld):

so, (0,7) (-4,9) and what else?

OpenStudy (triciaal):

try(5, -6)

OpenStudy (joftheworld):

its not a straight line..

OpenStudy (triciaal):

i don't know how to explain except 5 x - 7 is a multiple of 3

OpenStudy (joftheworld):

its like substituting (x,y)

OpenStudy (triciaal):

linear means line

OpenStudy (joftheworld):

yeah I know that .. do you have any clue how to do this ?

OpenStudy (triciaal):

didn't the ones i gave you work? has to do with the scale factor 3 did you read the post above 5 x - 7 is a multiple of 3 example when x = 5 5*5-7= 25-7 = 18 then divide by -3 to get y =-6 so the point was (5, -6)

OpenStudy (joftheworld):

the first one worked.. (-4,9)